OpenStudy (anonymous):

Please give an easy and fast way to solve this:

OpenStudy (anonymous):

\[x ^{2}-xc-2c ^{2}=0\]

OpenStudy (anonymous):

Use the quadratic formula.

OpenStudy (anonymous):

\[(a+b^{2}=a ^{2} + 2ab + b ^{2}\]

OpenStudy (anonymous):

solve what?

OpenStudy (anonymous):

but isn't' a quadratic equation of the type: \[a ^{2}+2ab+b ^{2}=0\] OR \[a ^{2}-2ab+b ^{2}=0\] My equation has two -ve signs.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

You have to get x = 2c

OpenStudy (anonymous):

(x+a)*(x-b)

OpenStudy (anonymous):

Please show the steps.

OpenStudy (anonymous):

x^2−xc−2*c^2=0 (x+c)(x-2c)=0

OpenStudy (yuki):

\[x = {c \pm \sqrt{c^2 - 4*z*(-2c^2)} \over 2}\]

OpenStudy (yuki):

lol

OpenStudy (anonymous):

x^2-x c-2c^2=0 x=-c+-sqrt(c^2-4*1*2c^2)/2*1

OpenStudy (anonymous):

OK. Thank You.

OpenStudy (anonymous):

c can be treated as a constant

OpenStudy (anonymous):

\[x^{2}-xc-c ^{2}=x^{2}-2xc+xc-c ^{2}=(x-2c)(x+c)=0 so we have x=2c or x=-c\]